# Communicating the gravitational-wave discoveries of the LIGO–Virgo–KAGRA Collaboration

### v2

@@ Acknowledgements
This material is based upon work supported by LIGO Laboratory which is a major facility fully funded by the US National Science Foundation and by the European Gravitational Observatory (EGO) which is funded by Centre National de la Recherche Scientifique (CNRS) in France, the Istituto Nazionale di Fisica Nucleare (INFN) in Italy, and the National Institute for Subatomic Physics (Nikhef) in the Netherlands. This paper has been assigned LIGO DCC number P2400039 (dcc.ligo.org/P2400039).
The Tactile Universe project was jointly funded by the University of Portsmouth and an STFC Public Engagement Legacy Award [ST/V001515/1]. CPLB acknowledges support of STFC grant ST/V005634/1. IC-C acknowledges support by the Spanish Agencia Estatal de Investigación through the Grant No. PID2021-125485NB-C21 funded by the MCIN/AEI/10.13039/501100011033, by the Generalitat Valenciana (Prometeo program for excellent research groups) through the Grant No. CIPROM/2022/49 and the Astrophysics and High Energy Physics program Grant No. ASFAE/2022/003 funded by MCIN and the European Union NextGenerationEU (PRTR-C17.I1). HM is supported by the UK Space Agency Grant Nos. ST/Y004922/1; ST/V002813/1; ST/X002071/1.

### v3

@@ Acknowledgements
We thank Jess McIver, the listeners of Coffee Break: Señal y Ruido, and the anonymous referees for constructive feedback on this manuscript. This material is based upon work supported by LIGO Laboratory which is a major facility fully funded by the US National Science Foundation and by the European Gravitational Observatory (EGO) which is funded by Centre National de la Recherche Scientifique (CNRS) in France, the Istituto Nazionale di Fisica Nucleare (INFN) in Italy, and the National Institute for Subatomic Physics (Nikhef) in the Netherlands This paper has been assigned LIGO DCC number P2400039 (dcc.ligo.org/P2400039/public).
The Tactile Universe project was jointly funded by the University of Portsmouth and an STFC Public Engagement Legacy Award [ST/V001515/1]. HM is supported by the UK Space Agency Grant Nos. ST/Y004922/1; ST/V002813/1; ST/X002071/1. CPLB acknowledges support of STFC grant ST/V005634/1. IC-C acknowledges support by the Spanish Agencia Estatal de Investigación through the Grant No. PID2021-125485NB-C21 funded by the MCIN/AEI/10.13039/501100011033, by the Generalitat Valenciana (Prometeo program for excellent research groups) through the Grant No. CIPROM/2022/49 and the Astrophysics and High Energy Physics program Grant No. ASFAE/2022/003 funded by MCIN and the European Union NextGenerationEU (PRTR-C17.I1), and by the European Horizon Europe staff exchange (SE) programme HORIZON-MSCA-2021-SE-01 Grant No. NewFunFiCO-101086251. ZD is grateful for support from the CIERA Board of Visitors Research Assistant Professorship. ACG is supported by the Dutch Research Council (NWO) with project number VI.Veni.212.047. DEH is supported by NSF grants AST-2006645 and PHY-2110507, as well as by the Kavli Institute for Cosmological Physics through an endowment from the Kavli Foundation. DK was supported by the Spanish Agencia Estatal de Investigación grants CNS2022-135440 and PID2022-138626NB-I00, funded by MICIU/AEI/10.13039/501100011033 and the ERDF/EU; and the Comunitat Autònoma de les Illes Balears through the Servei de Recerca i Desenvolupament and the Conselleria d’Educació i Universitats with funds from the Tourist Stay Tax Law (PDR2020/11 - ITS2017-006), from the European Union - NextGenerationEU/PRTR-C17.I1 (SINCO2022/6719) and from the European Union - European Regional Development Fund (ERDF) (SINCO2022/18146).